# The Missing Parts: Augmenting Fact Verification with Half Truth Detection

### v1

@@ Abstract
Fact verification systems typically assess whether a claim is supported by retrieved evidence, assuming that truthfulness depends solely on what is stated. However, many real-world claims are half-truths, factually correct yet misleading due to the omission of critical context. Existing models struggle with such cases, as they are not designed to reason about what is left unsaid. We introduce the task of half-truth detection, and propose PolitiFact-Hidden, a new benchmark with 15k political claims annotated with sentence-level evidence alignment and inferred claim intent. To address this challenge, we present TRACER, a modular re-assessment framework that identifies omission-based misinformation by aligning evidence, inferring implied intent, and estimating the causal impact of hidden content. TRACER can be integrated into existing fact-checking pipelines and consistently improves performance across multiple strong baselines. Notably, it boosts Half-True classification F1 by up to 16 points, highlighting the importance of modeling omissions for trustworthy fact verification. We will release both the benchmark and code upon publication.
@@
Yixuan Tang   Jincheng Wang   Anthony K.H. Tung Department of Computer Science, National University of Singapore yixuan@comp.nus.edu.sg, bertrand.wongjc@gmail.com, atung@comp.nus.edu.sg
@@ 1 Introduction
However, many real-world claims are not outright false but are still misleading due to the omission of critical context. These are often referred to as half-truths: statements that are factually correct but strategically incomplete Singamsetty et al. (2023a); Jaradat et al. (2024). Consider the example in Table 1, where a politician claims that unemployment has reached a 50-year low. While this statistic is factually accurate, it omits key information, such as the rise in part-time gig jobs and stagnant labor force participation, that undermines the implied narrative of broad economic success. Standard fact verification (FV) models, which focus on validating surface-level factuality, label such claims as true, failing to capture the misleading nature of selective omission.
@@ 2 Related Work
Omission-based misinformation, including half-truths, has received increasing attention. Singamsetty et al. (2023b) introduce controlled claim editing to expose omitted content, and Chen et al. (2022) propose generating implicit questions to recover missing context. Other datasets have incorporated related annotations, such as Cherry-picking Schlichtkrull et al. (2023) and Mixture Yang et al. (2022), which primarily capture conflicting evidence rather than omissions per se. These schemes focus on factual inconsistency (i.e., presence of both supporting and refuting evidence), rather than semantic incompleteness or intent-driven distortion. In contrast, our work targets half-truths, claims that are factually accurate but strategically omit Critical Hidden Evidence (CHE) that significantly alters interpretation. Closely related are efforts that explore the role of intent in misinformation, such as distinguishing disinformation through concealment and overstatement Rodríguez-Ferrándiz (2023); Lee and Lee (2024). We go beyond these by explicitly modeling the causal impact of hidden evidence on inferred intent without altering the original claim.
@@
Recent methods incorporate structured reasoning to improve factuality assessment. Program-guided models such as QACheck and ProgramFC Pan et al. (2023b) generate intermediate steps to support verification. Argumentation-based approaches, such as CHECKWHY Si et al. (2024), model causal links within evidence chains. Meanwhile, prompting-based methods like HiSS Zhang and Gao (2023a) and Flan-T5 Chung et al. (2022a) leverage large language models for step-by-step verification. Other work explores intent modeling using contrastive learning Yang et al. (2024) or refined retrieval Wang et al. (2024). Our work complements these efforts by introducing omission-aware reasoning and providing a modular framework that can be integrated into both structured and generative pipelines.
@@ 3 Task Formulation
Half-True Mostly True, Half True
@@
The task consists of two sub-problems:
1. Evidence Alignment: For each $e_{i}\in E$, classify whether it belongs to PE or HE. This is framed as a binary classification problem.
2. Causality Analysis: Among the hidden evidence, identify whether any subset constitutes CHE by evaluating its impact on the intent of the claim.
@@ 4 Dataset: PolitiFact-Hidden
Figure 1: Illustration of the annotation pipeline.
As illustrated in Figure 1, We develop a semi-automated annotation pipeline (Figure 1) combining GPT-4o-mini prompting and model-assisted refinement to label each claim with evidence alignment and intent metadata.
Evaluation Metric Requirement LLM-Human Agreement
Plausibility The inferred intent must not contradict the claim. 96.0%
Implicity The intent should be implied, not overtly stated. 96.0%
Sufficiency The description must be specific and informative. 91.0%
Readability The intent must be clearly and fluently expressed. 86.0%
@@ 4.1 Data Source and Label Schema
The dataset is built upon fact-checking articles from PolitiFact, which include both a concise claim and an accompanying verdict article. Unlike many other fact-checking sources, PolitiFact explicitly considers completeness in its rating criteria: a claim rated True must be both accurate and complete, while Mostly True and Half True indicate factual correctness with missing context (Holan, 2018). In contrast, Mostly False reflects the presence of conflicting evidences.
@@ 4.2 Annotation Pipeline
1. Relevance Check: Filter out irrelevant content using LLM-based entailment prompting.
2. Presentation Check: Assess whether the content is explicitly or implicitly stated in the claim.
3. Similarity Refinement: Use cosine similarity with XLM-RoBERTa embeddingsNils Reimers (2019) to refine edge cases and mitigate hallucinations.
Evidence is labeled as either:
  • Presented Evidence (PE): Explicitly or implicitly conveyed by the claim.
  • Hidden Evidence (HE): Relevant but not stated in the claim.
Manual inspection of 50 samples shows an 88% agreement between LLM predictions and human judgments, validating the alignment process.
@@
We define intent as the implied message the claim seeks to deliver. To extract intent:
1. Ruling Enhancement: Enhance ruling text by adding supporting evidence for clarity.
2. Intent Extraction: Use instruction-tuned prompting to extract the claim’s intended conclusion.
3. Quality Filtering: Filter extracted intents using four criteria, namely plausibility, implicity, sufficiency and readability.
The high agreement scores in Table 4 indicate that our LLM-assisted approach reliably captures high-quality intent for downstream reasoning.
@@ 5.2 Intent Generation
Factual claims often carry implicit conclusions shaped by omission. We define intent as the implied message or judgment a claim aims to convey. Understanding this latent intent is essential for determining whether omitted content is misleading.
To infer intent, we prompt-tune an LLM using input that includes the claim and its associated evidence context. This prompt-based formulation encourages the model to extract implicit conclusions without relying on manually predefined templates. The resulting intents serve as semantic anchors for later causal estimation.
We ensure that extracted intents meet four criteria, plausibility, implicity, sufficiency, and readability, as described in Section 4. Only high-quality intents are retained for downstream reasoning.
@@ 5.3 Causality Analysis
Not all hidden evidence is misleading. To distinguish Critical Hidden Evidence (CHE) from neutral omissions, we estimate the causal influence of each HE sentence on the inferred intent.
Inspired by abductive reasoning frameworks Chen et al. (2022), we generate candidate assumptions $A_{i}$ that must hold for the intent $Z$ to be valid. These assumptions are derived from hidden evidence through binary question generation and abstraction.
We then evaluate the impact of each $A_{i}$ using counterfactual prompting: given $do(A_{i}=\neg A_{i})$, does the intent $Z$ still hold? If not, $A_{i}$ is marked as causally important. For each validated assumption, we retrieve corresponding CHE from the HE pool by selecting sentences that either support or contradict it, based on semantic similarity and an NLI model that verifies logical entailment.
@@ 6.2 Intent Generation
Table 6: Intent generation performance.
@@
   Improvement $\uparrow$ 0.7 $\uparrow$ 2.2 $\uparrow$ 0.2 $\uparrow$ 3.8
@@
   Improvement $\uparrow$ 1.8 $\uparrow$ 3.6 $\uparrow$ 4.1 $\uparrow$ 6.4
@@ 6.3 Baselines
  • HiSS Zhang and Gao (2023b): a state-of-the-art instruction-following verifier that employs structured reasoning by decomposing the claim into multiple verifiable subclaims and evaluating them step by step.
@@
  • QACheck Pan et al. (2023a) and ProgramFC Pan et al. (2023c): agent-based fact-checkers. QACheck decomposes claims into sub-questions and verifies them with evidence. ProgramFC treats verification as a structured program of sub-tasks generated via in-context learning and executed by modular agents.
  • CHECKWHY Si et al. (2024) and Flan-T5 Chung et al. (2022b): prompting-based LLMs. CHECKWHY models causal reasoning through argument structures. Flan-T5 is identified as a strong fact verifier in hallucination evaluations.
@@ 6.4 Main Results
Table 7 shows that TRACER improves both accuracy and macro-F1 when added to strong reasoning-based backbones. For example, integrating TRACER with HiSS improves test accuracy from 78.3% to 81.9%, and macro-F1 from 59.4 to 65.7. Similarly, CoT benefits from TRACER with a 2.2 point gain in test accuracy and a 3.8-point increase in macro-F1. These gains are observed across both dev and test sets, indicating the robustness of TRACER as a general-purpose re-assessment module.
@@ 6.5 Qualitative Analysis
Claim: “Under the Obama economy, utility bills are higher.” This claim was labeled True by HiSS, as it aligns with data showing an increase in utility costs during President Obama’s term. However, TRACER inferred an intent to attribute blame for rising prices to Obama’s economic policies. It then retrieved omitted context showing that electricity prices rose even faster under the previous administration and followed a similar pattern across presidencies. This weakened the implied causal attribution and led TRACER to revise the label to Half-True.
@@ 7 Conclusion
This work introduces the task of half-truth detection, addressing claims that are factually correct but misleading due to omitted context. To support this, we introduce PolitiFact-Hidden, a new benchmark with annotated evidence alignment and intent. We propose TRACER, a novel framework that detects omission-based misinformation via intent modeling and causal reasoning over hidden content. Integrated with existing fact verification models, TRACER consistently improves performance, especially on Half-True cases, demonstrating the importance of reasoning about what is not said. This work highlights omission-aware verification as a critical next step for building trustworthy fact-checking systems, and establishes TRACER as a generalizable framework for tackling this underexplored but essential challenge.
@@
While TRACER demonstrates strong performance in identifying omission-based misinformation, several limitations remain. First, Our evaluation focuses on political discourse, as PolitiFact-Hidden is constructed from the PolitiFact corpus. While TRACER is designed to be model-agnostic and domain-independent, its effectiveness in other domains, such as health or finance, remains to be validated, especially where omission patterns may differ. Second, TRACER assumes that each claim expresses a coherent and inferable intent. However, real-world claims may be vague, ambiguous, or convey multiple overlapping intents, which can introduce noise in downstream reasoning. Future work may explore more robust modeling of claim pragmatics and intent uncertainty to extend TRACER’s applicability to broader, less structured scenarios.
Appendix A Prompts to Construct Politifact-Hidden
The following illustrates the exact prompt templates used to construct Politifact-Hidden, including annotation and evaluation.
Prompt:
You are tasked to determine the relevance of an evidence to a event.
You will be given a claim, the fact checking justification of this claim, and an evidence. Is the evidence irrelevant to the event?
Irrelevant: The evidence does not talk about one aspect of the event.
Relevant: The evidence talks about one aspect of the event even if it does not directly address the claim or shares the general topics of the event or simply reference to the original claim.
You do not need to focus on does the evidence support or refute the claim.
Evidence: {evidence}
@@ Appendix A Prompts to Construct Politifact-Hidden
Is the evidence relevant to the event?
A. Yes
B. No
Answer only one letter:
Figure 5: Prompt to determine evidence relevance to an event.
Prompt:
You are tasked to determine whether the evidence is presented in a claim.
You will be given a claim and an evidence. Is the evidence presented in the claim?
Presented should satisfy the following conditions:
  1. 1.
    The evidence partly or fully supports the claim. No contradiction is found.
  2. 2.
    The evidence supports the claim without further reasoning, because information is directly and explicitly stated in the claim.
Evidence: {evidence}
@@ A. Yes
Is the evidence presented in the claim?
A. Yes
B. No
Answer only one letter:
Figure 6: Prompt to determine whether evidence is explicitly presented in a claim.
Prompt:
You will be provided with the ruling and evidence from a fact-checking article. Your task is to enhance the clarity and depth of the ruling.
Definitions:
  • Ruling: A concise summary of the fact-checking article that includes the veracity rating of the claim.
  • Evidence: The supporting details and collected data related to the claim.
Requirements:
  • Identify Ambiguities: Review the ruling and evidence to pinpoint any unclear or incomplete information in the ruling.
  • Enrich with Evidence: Refer to the relevant parts of the evidence to expand the ruling. Ensure the enriched ruling explicitly explains how the evidence supports or contradicts the claim and connects directly to its veracity rating.
  • Create a Comprehensive Ruling: The enhanced ruling should independently present the full context of the fact-checking process and the rationale for the given rating.
Evidence: {evidence}
@@
Figure 7: Prompt to enrich fact-checking ruling using the given evidence.
Prompt:
A claim would convey implicit intents. You are required to determine the intent of a claim based on context in Ruling.
Definition:
  • Claim: The claim that is checked.
  • Ruling: Text to determine veracity and explain how the claim would shape people’s understanding.
  • Intent: The understanding of the event that the speaker wants to shape, which is not directly presented in the claim.
(3 Examples are omitted)
Requirements:
  1. 1.
    Intent must be checkable. For example, "people should do something" is not checkable because it does not happen until now.
  2. 2.
    Output intent in <>.
  3. 3.
    Please think step by step. First write your rationale, then the intent.
Claim: {claim}
@@
Figure 8: Prompt to extract intent of a claim from a enhanced fact-checking ruling.
Prompt:
You are required to determine whether the intended conclusion is a plausible intent of the claim.
Please rate using the following scale:
  • 0 (not plausible): The claim contradicts the intended conclusion.
  • 1 (plausible): The claim does not contradict the intended conclusion.
Claim: {claim}
@@
Figure 9: Prompt to evaluate the plausibility of an intended conclusion based on a claim.
Prompt:
You are required to determine whether the intended conclusion conveys the implicit meaning of the claim.
Choose from {0, 1}:
  • 0 (not implicit): The intended conclusion simply rephrases some part of the claim. It does not convey any implicit meaning of the claim.
  • 1 (implicit): The intended conclusion reveals implicit information that is not explicitly stated in the claim.
Claim: {claim}
Intended Conclusion: {intent}
Output only one digit.
Figure 10: Prompt to evaluate whether the intended conclusion expresses the implicit meaning of the claim.
Prompt:
You are required to determine whether the intended conclusion is sufficient, meaning that it is understandable within the scope of general knowledge.
Choose from {0, 1}:
  • 0 (not sufficient): The intended conclusion has obvious ambiguous references and is not understandable. For example, it uses unclear terms like “the claim”.
  • 1 (sufficient): The intended conclusion is clearly referenced and understandable on its own.
Intended Conclusion: {intent}
Output only one digit.
Figure 11: Prompt to evaluate whether the intended conclusion is sufficiently clear and self-contained.
Prompt:
You are required to determine whether the intended conclusion is readable.
Choose from {0, 1}:
  • 0 (not readable): The intended conclusion is not readable and is overly complicated.
  • 1 (readable): The intended conclusion is readable and understandable.
Intended Conclusion: {intent}
Output only one digit.
Figure 12: Prompt to evaluate whether the intended conclusion is readable.
Appendix B Prompts Used in TRACER
The following illustrates the exact prompt templates used at each stage of the TRACER framework, including implicit question generation, assumption inference, causality evaluation, and final re-assessment.
Prompt:
A claim can be literally accurate but still misleading in an implicit way.
Your task is to identify the important implicit questions addressed by the evidence.
Steps:
  1. 1.
    Read the evidence below carefully to understand the full context and the topics it covers.
  2. 2.
    Assume the claim is true. What important implicit yes-no questions should be asked to verify the intended conclusion, rather than just the literal accuracy of the claim?
  3. 3.
    Generate 1–3 such implicit questions.
  4. 4.
    Each question should be enclosed in its own angle brackets <>.
  5. 5.
    All implicit questions must be yes-no questions.
(Examples are omitted.)
Claim: {claim}
Intended conclusion: {intent}
Evidence:
{evidence}
Figure 13: Prompt to generate implicit questions.
Prompt:
A claim could be literally accurate but still misleading because of its intended conclusion.
Your task is to determine what assumptions the intended conclusion is based on, besides the claim.
Definition:
  • Claim: A statement assumed to be true.
  • Intended conclusion: The intended conclusion of the claim, which needs checking.
  • Questions: Some important questions when checking the claim.
  • Assumptions: The assumptions that the intended conclusion is based on, besides the claim.
Steps:
  1. 1.
    Read the claim, intended conclusion, and questions.
  2. 2.
    Assuming the claim is correct, what assumptions does the question imply should serve as the basis for the intended conclusion?
  3. 3.
    Output a 1–3 sentence rationale, followed by 1–{assumption_max_number} assumptions.
    Each assumption should be enclosed in angle brackets <> and separated by ||.
Requirements:
  1. 1.
    Ensure that each assumption can independently convey its meaning.

### v2

@@ Abstract
Fact verification systems typically assess whether a claim is supported by retrieved evidence, assuming that truthfulness depends solely on what is stated. However, many real-world claims are half-truths, factually correct yet misleading due to the omission of critical context. Existing models struggle with such cases, as they are not designed to reason about omitted information. We introduce the task of half-truth detection, and propose PolitiFact-Hidden, a new benchmark with 15k political claims annotated with sentence-level evidence alignment and inferred claim intent. To address this challenge, we present TRACER, a modular re-assessment framework that identifies omission-based misinformation by aligning evidence, inferring implied intent, and estimating the causal impact of hidden content. TRACER can be integrated into existing fact-checking pipelines and consistently improves performance across multiple strong baselines. Notably, it boosts Half-True classification F1 by up to 16 points, highlighting the importance of modeling omissions for trustworthy fact verification. The benchmark and code are available via https://github.com/tangyixuan/TRACER.
@@
Yixuan Tang    Jincheng Wang    Anthony K.H. Tung School of Computing, National University of Singapore yixuan@comp.nus.edu.sg, bertrand.wongjc@gmail.com, atung@comp.nus.edu.sg
@@ 1 Introduction
However, many real-world claims are not outright false but are still misleading due to the omission of critical context. Misinformation can evolve dynamically when propagated under different political stances  Chong et al. (2025), these are often referred to as half-truths, i.e. statements that are factually correct but strategically incomplete Singamsetty et al. (2023); Jaradat et al. (2024). Consider the example in Table 1, where a politician claims that unemployment has reached a 50-year low. While this statistic is factually accurate, it omits key information, such as the rise in part-time gig jobs and stagnant labor force participation, that undermines the implied narrative of broad economic success. Standard fact verification (FV) models, which focus on validating surface-level factuality, label such claims as true, failing to capture the misleading nature of selective omission.
@@ 2 Related Work
Omission-based misinformation, including half-truths, has received increasing attention. Singamsetty et al. (2023) introduce controlled claim editing to expose omitted content, and Chen et al. (2022) propose generating implicit questions to recover missing context. Other datasets have incorporated related annotations, such as Cherry-picking Schlichtkrull et al. (2023) and Mixture Yang et al. (2022), which primarily capture conflicting evidence rather than omissions per se. These schemes focus on factual inconsistency (i.e., presence of both supporting and refuting evidence), rather than semantic incompleteness or intent-driven distortion. In contrast, our work targets half-truths, claims that are factually accurate but strategically omit Critical Hidden Evidence (CHE) that significantly alters interpretation. Closely related are efforts that explore the role of intent in misinformation, such as distinguishing disinformation through concealment and overstatement Rodríguez-Ferrándiz (2023); Lee and Lee (2024). Tang et al. (2025) uncover the comprehensive view of events by mitigating selective presentation of information, they do not integrate downstream fact verification. We go beyond these by explicitly modeling the causal impact of Hidden Evidence on inferred intent without altering the original claim.
@@
Recent methods incorporate structured reasoning to improve factuality assessment. Program-guided models such as QACheck and ProgramFC Pan et al. (2023b) generate intermediate steps to support verification Tang et al. (2021). Argumentation-based approaches, such as CHECKWHY Si et al. (2024), model causal links within evidence chains. Meanwhile, prompting-based methods like HiSS Zhang and Gao (2023) and Flan-T5 Chung et al. (2022) leverage large language models for step-by-step verification. Other work explores intent modeling using contrastive learning Yang et al. (2024) or refined retrieval Wang et al. (2024). Our work complements these efforts by introducing omission-aware reasoning and providing a modular framework that can be integrated into both structured and generative pipelines.
@@ 3 Task Formulation
Half-True Mostly True, Half-True
@@ 4 Dataset: PolitiFact-Hidden
Figure 2: Illustration of the semi-automated annotation pipeline for constructing PolitiFact-HIDDEN, combining GPT-4o-mini prompting with human quality control.
As illustrated in Figure 2, we develop a semi-automated annotation pipeline (Figure 2) combining GPT-4o-mini prompting and model-assisted refinement to label each claim with evidence alignment and Intent.
Dimension Requirement LLM-Positive Human Confirmed Agreement
Plausibility The inferred intent must not contradict the claim. 95 94 98.9%
Implicity The intent should be implied, not overtly stated. 94 93 98.9%
Sufficiency The description must be specific and informative. 81 80 98.8%
Readability The intent must be clearly and fluently expressed. 76 70 92.1%
@@ 4.1 Data Source and Label Schema
The dataset is built upon fact-checking articles from PolitiFact, which include both a concise claim and an accompanying verdict article. Unlike many other fact-checking sources, PolitiFact explicitly considers completeness in its rating criteria: a claim rated True must be both accurate and complete, while Mostly True and Half-True indicate factual correctness with missing context (Holan, 2018). In contrast, Mostly False reflects the presence of conflicting evidences.
@@ 4.2 Annotation Pipeline
  1. 1.
    Relevance Check: Filter out irrelevant content using LLM-based entailment prompting.
  2. 2.
    Presentation Check: Assess whether the content is explicitly or implicitly stated in the claim.
  3. 3.
    Similarity Refinement: Use cosine similarity with XLM-RoBERTa embeddingsNils Reimers (2019) to refine edge cases and mitigate hallucinations.
Evidence is labeled as either PE or HE. Manual inspection of 50 samples showed an 88% agreement between LLM predictions and human judgments, validating the alignment process.
@@
A key element of half-truth detection is the claim’s Intent, i.e., the implied message or judgment it seeks to convey. Intents are extracted in 3 steps:
  1. 1.
    Ruling Enhancement: Enhance ruling text by adding supporting evidence for clarity.
  2. 2.
    Intent Extraction: Use instruction-tuned prompting to extract the claim’s intended conclusion.
  3. 3.
    Quality Filtering: Filter extracted intents using four criteria, namely plausibility, implicity, sufficiency and readability.
To validate the quality of LLM-based filtering, we had two human annotators independently assess 100 samples across the same four evaluation dimensions. Agreement between the LLM and both annotators was high (92.1-98.9% across dimensions), suggesting that the LLM-assisted approach reliably captures high-quality intents for downstream reasoning. The full intent evaluation prompts are provided in Appendix A.
@@ 5.2 Intent Generation
Understanding this latent intent is essential for determining whether omitted content is misleading. As described in Section 4.2, we prompt-tune an LLM using input that includes the claim and its associated evidence context to infer intent. This prompt-based formulation encourages the model to extract implicit conclusions without relying on manually predefined templates. The resulting intents serve as semantic anchors for subsequent causality analysis.
@@ 5.3 Causality Analysis
While assumptions are derived from HE, not all HE sentences directly affect the plausibility of the intent. Many are tangential or neutral. To distinguish Critical Hidden Evidence (CHE) from neutral omissions, we estimate the causal influence of each HE sentence on the inferred intent.
Inspired by abductive reasoning frameworks Chen et al. (2022), we generate candidate assumptions $A_{i}$ that must hold for the intent $Z$ to be valid. These assumptions are derived from evidence through binary question generation and abstraction.
We then evaluate the impact of each $A_{i}$ using counterfactual prompting: given $do(A_{i}=\neg A_{i})$, does the intent $Z$ still hold? If not, $A_{i}$ is marked as causally important. For each validated assumption, we retrieve corresponding CHE from the HE pool by selecting sentences that either support or contradict it, based on semantic similarity and an NLI model that verifies logical entailment. This two-step refinement prevents irrelevant or weakly related evidence from being misclassified as CHE.
@@ 6.2 Intent Generation
Table 6: Performance of intent generation methods.
@@
   Improvement $\uparrow$ 0.7 $\uparrow$ 2.2 $\uparrow$ 0.2 $\uparrow$ 3.7
@@
   Improvement $\uparrow$ 1.8 $\uparrow$ 3.6 $\uparrow$ 4.1 $\uparrow$ 6.3
@@ 6.3 Baselines
  • HiSS Zhang and Gao (2023): a state-of-the-art instruction-following verifier that employs structured reasoning by decomposing the claim into multiple verifiable subclaims and evaluating them step by step.
@@
  • QACheck Pan et al. (2023a) and ProgramFC Pan et al. (2023b): agent-based fact-checkers. QACheck decomposes claims into sub-questions and verifies them with evidence. ProgramFC treats verification as a structured program of sub-tasks generated via in-context learning and executed by modular agents.
  • CHECKWHY Si et al. (2024) and Flan-T5 Chung et al. (2022): prompting-based LLMs. CHECKWHY models causal reasoning through argument structures. Flan-T5 is identified as a strong fact verifier in hallucination evaluations.
To ensure fairness, we evaluate all baselines using GPT-4o-mini, except in cases where prior work demonstrates that a different backbone yields stronger performance. For HiSS, we find GPT-3.5-turbo consistently outperforms GPT-4o-mini.
@@ 6.4 Main Results
Table 7 shows that TRACER improves both accuracy and macro-F1 when added to strong reasoning-based backbones. For example, integrating TRACER with HiSS improves test accuracy from 78.3% to 81.9%, and macro-F1 from 59.4 to 65.7. Similarly, CoT benefits from TRACER with a 2.2 point gain in test accuracy and a 3.7-point increase in macro-F1. These gains are observed across both dev and test sets, indicating the robustness of TRACER as a general-purpose re-assessment module.
@@
Per-Class Performance.
To further examine TRACER’s effect on fact verification, we report per-class performance for the top-performing models. As shown in Table 9, TRACER substantially improves the classification of Half-True claims while also maintaining or slightly enhancing performance on True and False claims. This confirms that the observed gains are not achieved at the expense of other classes.
Generalization.
To examine the generalization of TRACER, we evaluate it with the open-source LLaMA2-7B model as the base verifier on the top-performing HiSS framework. With TRACER, accuracy improves from 78.2 to 82.3 and Macro-F1 from 59.1 to 65.4. A breakdown of per-class performance and a follow-up analysis of results over different claim lengths is provided in Appendix C.
Method True Half-True False
CoT 52.9 52.8 87.1
CoT +RA 56.7 $\uparrow$ 3.8 60.2 $\uparrow$ 7.4 87.1 (–)
HiSS 44.7 44.4 88.9
HiSS +RA 46.6 $\uparrow$ 1.9 60.5 $\uparrow$ 16.1 90.1 $\uparrow$ 1.2
Table 9: Per-class F1 scores on the test set.
@@ 6.5 Qualitative Analysis
Claim: “Under the Obama economy, utility bills are higher.” This claim was labeled True by HiSS, as it aligns with data showing an increase in utility costs during President Obama’s term. However, TRACER inferred an intent to attribute blame for rising prices to Obama’s economic policies. It then retrieved CHE showing that electricity prices rose even faster under the previous administration and followed a similar pattern across presidencies. This weakened the implied causal attribution and led TRACER to revise the label to Half-True.
@@ 7 Conclusion
This work introduces the task of half-truth detection, addressing claims that are factually correct but misleading due to omitted context. To support this, we introduce PolitiFact-Hidden, a new benchmark with annotated evidence alignment and intent. We propose TRACER, a novel framework that detects omission-based misinformation via intent modeling and causal reasoning over hidden content. Integrated with existing fact verification models, TRACER consistently improves performance, especially on half-truths, demonstrating the importance of reasoning about omitted information. This work highlights omission-aware verification as a critical next step for building trustworthy fact-checking systems, and establishes TRACER as a generalizable framework for tackling this underexplored but essential challenge.
@@
While TRACER demonstrates strong performance in identifying omission-based misinformation, several limitations remain. First, our evaluation focuses on political discourse, as PolitiFact-Hidden is constructed from the PolitiFact corpus. While TRACER is designed to be model-agnostic and domain-independent, its effectiveness in other domains, such as health or finance, remains to be validated, especially where omission patterns may differ. Second, TRACER assumes that each claim expresses a coherent and inferable intent. However, real-world claims may be vague, ambiguous, or convey multiple overlapping intents, which can introduce noise in downstream reasoning. Future work may explore more robust modeling of claim pragmatics and intent uncertainty to extend TRACER’s applicability to broader scenarios.
Acknowledgments
This research is supported by the Ministry of Education, Singapore, under its MOE AcRF TIER 3 Grant (MOE-MOET32022-0001).
Appendix A Prompts for Constructing Politifact-Hidden
This section presents the prompt templates employed in building the Politifact-Hidden dataset.
Prompt: Evidence Relevance Classification You are tasked to determine the relevance of an evidence to an event. You will be given a claim, the fact-checking justification of this claim, and an evidence. Is the evidence irrelevant to the event? Irrelevant: The evidence does not talk about one aspect of the event. Relevant: The evidence talks about one aspect of the event even if it does not directly address the claim or shares the general topics of the event or simply reference to the original claim. You do not need to focus on does the evidence support or refute the claim. Evidence: {evidence}
@@ Appendix A Prompts for Constructing Politifact-Hidden
Is the evidence relevant to the event? A. Yes B. No Answer only one letter:
Prompt: Evidence Presence Classification You are tasked to determine whether the evidence is presented in a claim. You will be given a claim and evidence. Is the evidence presented in the claim? Presented should satisfy the following conditions: 1. The evidence partly or fully supports the claim. No contradiction is found. 2. The evidence supports the claim without further reasoning, because information is directly and explicitly stated in the claim. Evidence: {evidence}
@@
Is the evidence presented in the claim? A. Yes B. No Answer only one letter:
Prompt: Enrich Fact-Checking Ruling with Evidence Given. You will be provided with the ruling and evidence from a fact-checking article. Your task is to enhance the clarity and depth of the ruling. Definitions: Ruling: A concise summary of the fact-checking article that includes the veracity rating of the claim. Evidence: The supporting details and collected data related to the claim. Requirements: Identify Ambiguities: Review the ruling and evidence to pinpoint any unclear or incomplete information in the ruling. Enrich with Evidence: Refer to the relevant parts of the evidence to expand the ruling. Ensure the enriched ruling explicitly explains how the evidence supports or contradicts the claim and connects directly to its veracity rating. Create a Comprehensive Ruling: The enhanced ruling should independently present the full context of the fact-checking process and the rationale for the given rating. Evidence: {evidence}
@@
Prompt: Intent Analysis A claim would convey implicit intents. You are required to determine the intent of a claim based on context in Ruling. Definition: Claim: The claim that is checked. Ruling: Text to determine veracity and explain how the claim would shape people’s understanding. Intent: The understanding of the event that the speaker wants to shape, which is not directly presented in the claim. (3 Examples are omitted) Requirements: 1. Intent must be checkable. For example, "people should do something" is not checkable because it does not happen until now. 2. Output intent in <>. 3. Please think step by step. First write your rationale, then the intent. Claim: {claim}
@@
To avoid repetition, we use colors in the prompts to denote different evaluation dimensions, which are assessed independently in practice.
Generated Intent Evaluation (4 dimensions) You are required to determine whether the intended conclusion is a plausible intent of the claim / conveys the implicit meaning of the claim / readable / sufficient, meaning that it is understandable within the scope of general knowledge. Please rate using the following scale: 0 (not plausible): The claim contradicts the intended conclusion. 1 (plausible): The claim does not contradict the intended conclusion. 0 (not implicit): The intended conclusion simply rephrases some part of the claim. It does not convey any implicit meaning of the claim. 1 (implicit): The intended conclusion reveals implicit information that is not explicitly stated in the claim. 0 (not readable): The intended conclusion is not readable and is overly complicated. 1 (readable): The intended conclusion is readable and understandable. 0 (not sufficient): The intended conclusion has obvious ambiguous references and is not understandable. For example, it uses unclear terms like “the claim”. 1 (sufficient): The intended conclusion is clearly referenced and understandable on its own. Claim: {claim}
@@
Appendix B Prompt Templates Used in TRACER
This appendix provides the complete prompt templates employed at each stage of the TRACER framework. We include prompts for implicit question generation, assumption inference, causality evaluation, and final re-assessment.
Prompt: Implicit Questions Generation. A claim can be literally accurate but still misleading in an implicit way. Your task is to identify the important implicit questions addressed by the evidence. Steps: 1. Read the evidence below carefully to understand the full context and the topics it covers. 2. Assume the claim is true. What important implicit yes-no questions should be asked to verify the intended conclusion, rather than just the literal accuracy of the claim? 3. Generate 1–3 such implicit questions. 4. Each question should be enclosed in its own angle brackets <>. 5. All implicit questions must be yes-no questions. (Examples are omitted.) Claim: {claim} Intended conclusion: {intent} Evidence: {evidence}
Prompt: Assumption Generation A claim could be literally accurate but still misleading because of its intended conclusion. Your task is to determine what assumptions the intended conclusion is based on, besides the claim. Definition: Claim: A statement assumed to be true. Intended conclusion: The intended conclusion of the claim, which needs checking. Questions: Some important questions when checking the claim. Assumptions: The assumptions that the intended conclusion is based on, besides the claim. Steps: 1. Read the claim, intended conclusion, and questions. 2. Assuming the claim is correct, what assumptions does the question imply should serve as the basis for the intended conclusion? 3. Output a 1–3 sentence rationale, followed by 1–{assumption_max_number} assumptions.
Each assumption should be enclosed in angle brackets <> and separated by $|$.
Requirements: 1. Ensure that each assumption can independently convey its meaning.